# An End-to-End Model For Logits Based Large Language Models Watermarking

### v1

@@ Appendix C Effect on LLM Temperature
Figure 9: Detection performance on various LLM temperatures.
Fig. 9 presents the F1 scores of different watermarking methods across various LLM temperatures under clean and edited conditions, including synonymous substitution, copy-paste attack, and paraphrasing. The rightmost subfigure (e) reports the average F1 score across all conditions, highlighting overall robustness.
@@
Our method consistently outperforms all baselines, achieving an average F1 score of 0.976, which surpasses KGW and Unigram (both 0.930) by 4.9% and significantly outperforms Unbiased (0.841) and DiPmark (0.832) by 16.1% and 17.3%, respectively. Under clean conditions (a), all methods perform well, but our approach maintains a slight edge in stability across different temperatures. In synonymous substitution (b) and the copy-paste attack (c), our method demonstrates superior robustness, maintaining F1 scores near 1.0 while competing approaches experience noticeable declines. For paraphrasing (d), where detection is most challenging, our method exhibits a substantial advantage, outperforming the strongest baseline by over 7%. These results underscore the effectiveness of our method in maintaining high watermark detectability across varying LLM temperatures and adversarial conditions, outperforming traditional statistical approaches while ensuring robustness in real-world scenarios.
Appendix D Effect on LLM Sampling Strategy
@@
Table 7: Performance of the LLM watermark methods. CL: Clean watermark sample; SS: Synonymous substitution; CP: Copy-paste attack; PA; Paragraphing; PPL; text perplexity. The best is in bold, and the second-best is underlined.
Llama2-7B
@@
Appendix E Watermark Sample
@@
Figure 10 presents three pairs of non-watermarked and watermarked samples generated from the same prompt of our method. The tokens are colored based on the output values retrieved from the watermark encoder. The watermarked samples maintain high quality while featuring a greater proportion of green-list tokens compared to red-list tokens. This is achieved by increasing the probabilities of green-list tokens during the generation process.
Appendix F Model Training Details
@@
Learning rate 1-e4
Batch size 16
Training step 35000
@@
Appendix G Editing Configuration
@@
Paraphrasing Paraphrasing the text while maintaining the original meaning with deep paraphrase model Dipper(Krishna et al., 2023). Lex diversity: 60
@@
Appendix H Comparison with Post-Generation Watermarking Methods
Appendix I Statistically Significant Improvement
Table 10: Paired t-test results comparing our model with competitors. Mean Diff. represents the difference between our model and each competitor’s mean performance. All p-values below 0.05 indicate statistically significant improvements.
Llama2-7B
Method PA-Trials Mean ± Std Mean Diff. p-value
Method CP-Trials Mean ± Std Mean Diff. p-value
Method PPL-Trials Mean ± Std Mean Diff. p-value
To validate the statistical significance of our model’s improvements, we conducted robustness and quality experiments using a paired t-test. Specifically, we performed five trials to evaluate robustness against paraphrasing (PP) and copy-paste (CP) attacks, as well as the perplexity (PPL) of watermarked text. All trials were conducted using the same set of prompts, with multinomial sampling (same seed across methods in the same trial for consistency), resulting in diverse outcomes across the five trials. The watermark strength parameters ($\delta$) for KGW, Unigram, and our method were set to 2, 2, and 1.25, respectively, while the $k$-value for our method was set to 20. The results, shown in Table 10, indicate that our model achieves statistically significant improvements in both robustness and quality compared to the competitors. This is supported by p-values consistently below the 0.05 threshold.
Appendix J Watermark Efficiency
LLM Setting Generation Detection
Time (s) Memory (GB) Time (s) Memory (GB)
OPT-1.3B w/o WM 2.557 5.900 0.003 0.008
OPT-1.3B w/ WM 2.769 5.900 0.003 0.008
Llama2-7B w/o WM 5.204 15.793 0.005 0.008
Llama2-7B w/ WM 8.362 15.793 0.005 0.008
In Table11, we assess the computational time overhead and GPU memory usage of our method. For OPT-1.3B, the lightweight encoder design results in only an 8.3% increase in generation time for watermarked text, with no change in maximum GPU memory usage since the encoder is invoked after each token generation. For Llama2-7B, our method increases the generation time by 60.6%, mainly due to the embedding transformation from Llama to OPT, as the tokenizer cannot be accelerated by the GPU and is called at each step. The time overhead for watermark embedding can be mitigated through parallel tokenization, reducing the time complexity by up to $1/k$. In terms of watermark detection, our decoder operates efficiently, requiring only negligible time and memory consumption.
Appendix K Examples for online text modification
@@ Appendix I Statistically Significant Improvement
Copy-paste attack Please randomly copy the below given text and paste to the provided human-written text to form the new text: Climate change is one of the most pressing environmental issues of our time, posing significant threats to biodiversity worldwide. Global warming is a critical issue that affects various aspects of our planet. Climate change is one of the most pressing environmental issues of our time, posing significant threats to biodiversity worldwide. It leads to melting ice caps, rising sea levels, and more frequent natural disasters. These changes have profound impacts on both human societies and wildlife. …

### v2

@@ A.2 Effect on Sentence Length
Figure 8: Detection performance (F1↑) on watermark sentences with various length.
Fig. 8 presents the F1 score of various watermarking methods across different sentence lengths for three LLMs: OPT-1.3B, Llama2-7B, and Qwen2.5-7B. The results demonstrate that across all methods, F1 scores generally improve as sentence length increases, indicating enhanced watermark detectability in longer text. Our method consistently achieves strong performance across all models and sentence lengths. Notably, for shorter sentences (length $\leq 50$), our method attains higher F1 scores compared to other baselines, specially with an average improvement of 8% over DiPmark and Unbiased. For longer sentences (length $\geq 100$), all methods converge toward near-optimal F1 scores ($\approx 1.00$).
A.3 Effect on LLM Temperature
Figure 9: Detection performance (F1↑) on various LLM temperatures.
@@ A.3 Effect on LLM Temperature
Fig. 9 presents the F1 scores of different watermark methods across various LLM temperatures under clean and edited conditions, including synonymous substitution, copy-paste attack, and paraphrasing. The rightmost subfigure (e) reports the average F1 score across all conditions, highlighting overall robustness. Our method consistently outperforms all baselines, achieving an average F1 score of 0.976, which surpasses KGW and Unigram (both 0.930) by 4.9% and significantly outperforms Unbiased (0.841) and DiPmark (0.832) by 16.1% and 17.3%, respectively. Under clean conditions (a), all methods perform well, but our approach maintains a slight edge in stability across different temperatures. In synonymous substitution (b) and the copy-paste attack (c), our method demonstrates superior robustness, maintaining F1 scores near 1.0 while competing approaches experience noticeable declines. For paraphrasing (d), where detection is most challenging, our method exhibits a substantial advantage, outperforming the strongest baseline by over 7%. Lowering the temperature makes the LLM more deterministic, increasing the likelihood of sampling high-logit tokens. Since our model perturbs the top-logits, a lower temperature favors selecting tokens from our red/green lists over those in the grey list (see Section 3.2), thereby improving performance.
A.4 Effect on LLM Sampling Strategy
@@
Table 7: Detection performance (F1↑) of the LLM watermark methods with multinomial sampling and beam search. CL: Clean watermark sample; SS: Synonymous substitution; CP: Copy-paste attack; PA; Paragraphing. Detection Performance is evaluated on Llama2-7B. The best is in bold, and the second-best is underlined.
@@
A.5 False Positive Rate on Clean Watermark Sentence
Table 8: Best F1 FPR on clean watermark samples.
LLM Best F1 FPR↓
OPT-1.3B 0.00
Llama2-7B 0.03
Qwen2.5-7B 0.01
Mistral-7B 0.01
Llama3-8B 0.01
Llama3.2-3B 0.01
Avg. 0.01
We present the best F1 FPR for various LLMs of our method in Table 8. Notably, OPT-1.3B achieves an FPR of 0.00, Llama2-7B is at 0.03, and all other models (Qwen2.5-7B, Mistral-7B, Llama3-8B, and Llama3.2-3B) consistently achieve an FPR of 0.01. With an average FPR of just 0.01, our method demonstrates exceptionally low false positive rates across different LLMs.
A.6 Statistically Significant Improvement
Table 9: Paired t-test results comparing our model with competitors. Detection Performance is evaluated on Llama2-7B. PPL: Text perplexity with Llama2-13B. Mean Diff. represents the difference between our model and each competitor’s mean performance. All p-values below 0.05 indicate statistically significant improvements.
Method PA-Trials Mean ± Std Mean Diff. p-value
Method CP-Trials Mean ± Std Mean Diff. p-value
Method PPL-Trials Mean ± Std Mean Diff. p-value
To validate the statistical significance of our model’s improvements, we conducted robustness and quality experiments using a paired t-test in Table 9. Specifically, we performed five trials to evaluate robustness against paraphrasing (PP) and copy-paste (CP) attacks, as well as the perplexity (PPL) of watermarked text. All trials were conducted using the same set of prompts, with multinomial sampling (same seed across methods in the same trial for consistency), resulting in diverse outcomes across the five trials. The watermark strength parameters ($\delta$) for KGW, Unigram, and our method were set to 2, 2, and 1.25, respectively, while the $k$-value for our method was set to 20. The results, shown in Table 9, indicate that our model achieves statistically significant improvements in both robustness and quality compared to the competitors. This is supported by p-values consistently below the 0.05 threshold.
Appendix B Watermark Analysis
B.1 Watermark Sample
@@
Fig. 10 presents three pairs of non-watermarked and watermarked samples generated from the same prompt of our method. The tokens are colored based on the output values retrieved from the watermark encoder. The watermarked samples maintain high quality while featuring a greater proportion of green-list tokens compared to red-list tokens. The ratios between NWM/WM texts are not similar at all, the NWM texts have red/green token counts of 69:53, 61:65, and 74:75 (ratios 1.30, 0.94, and 0.99). In contrast, the WM texts show counts of 42:83, 50:99, and 53:79 (ratios 0.51, 0.51, and 0.67). This is achieved by increasing the probabilities of green-list tokens during the generation process.
B.2 Source of Robustness
We argue that robustness arises from two factors:
  1. 1.
    Watermark Decoder: By retrieving the red-green token lists with our watermark encoder (see Sec. 4.3), we can statistically compute token ratios that provide guarantees similar to provable p-values (analogous to KGW), which is critical for text forensics. As shown in Table 10, although our neural decoder—benefiting from end-to-end optimization with a noise layer—generally outperforms the statistical decoder, the latter remains competitive with strong baselines. At a fixed 1% FPR, the statistical decoder achieves a TPR of 0.96 in CP and a TPR of 0.65 in PA, compared to KGW’s 0.81 (CP) and 0.50 (PA) as well as Unigram’s 0.56 (CP) and 0.61 (PA). Since statistical decoder based solely on the red/green partition, and our neural decoder outperformed the statistical one, demonstrating that end-to-end training with a noise layer enhances robustness.
    Table 10: Detection performance of our neural decoder and statistical decoder.
    Method 1%FPR TPR↑ Best F1↑
    CL SS CP PA CL SS CP PA
    KGW ($\delta=2$) 1.00 0.94 0.81 0.50 1.00 0.96 0.92 0.82
    Unigram ($\delta=2$) 1.00 0.98 0.56 0.61 1.00 0.98 0.85 0.92
    Ours (Neural Decoder, $\delta=1.25$, $k=20$) 0.99 0.96 0.96 0.75 1.00 0.98 0.97 0.92
    Ours (Statistical Decoder, $\delta=1.25$, $k=20$) 0.95 0.91 0.96 0.65 0.97 0.96 0.97 0.85
  2. 2.
    Red/Green Partition: We measure the KL divergence (KLD) between token distributions of watermark (WM) and non-watermark (NWM) sentences (see Fig. 11) to evaluate the context independence (CI) of our partition. A purely CI partition (e.g., Unigram) shows high KLD due to token biases, while our method achieves a KLD of 0.12—about half of Unigram’s 0.21 and above KGW’s 0.03—indicating a balance between context-dependent (CD) and CI schemes. Our ablation study on context size after paraphrasing is shown in Table 11. Our adaptive partition lets the encoder use strong CD features when available and fall back to a CI approach when necessary, which is crucial for our robustness.
    Table 11: Detection Performance under different context size.
    Context Size 1%FPR TPR↑ (PA)
    2 0.69
    4 0.76
    8 0.79
    10 0.80
B.3 Risk of Watermark Stolen
Acquisition of Different Watermarks. It is NOT necessary to retrain the entire network from scratch for each user-specific watermark. The proposed end-to-end framework is flexible enough to inject key-driven randomness at multiple stages (input, model parameters, or output). In practice, one can fine-tune (FT) and apply key-conditioned post-processing without retraining. Different watermarks can be guaranteed by incorporating a key-driven bias logit $\mathbf{l}_{B}\in\{-1,1\}^{|\mathcal{V}|}$ (of vocabulary size) into the final watermark logits via $\hat{\mathbf{l}}=\mathbf{l}+\delta\cdot(\mathbf{l}_{W}+\mathbf{l}_{B})$, which $\mathbf{l}_{B}$ is only applied to the top-$k$ logits. Since $\mathbf{l}_{B}$ is derived from a unique key, statistical analysis demonstrates that each item in the clipped output remains unchanged with probability 0.5 and changes with probability 0.5 (assuming each $i$ item, $\mathbf{l}^{(i)}_{W}\sim\text{Uniform}(\{-1,1\})$, $\mathbf{l}^{(i)}_{B}\sim\text{Uniform}(\{-1,1\})$ and all items are independent). The number of mismatches between watermarks generated using different keys follows a binomial distribution $\text{Bin}(k,0.5)$, implying that the probability of obtaining identical watermarks is negligibly small. E.g. $k=20$, the probability of identical watermark is $0.5^{20}=9.54\times 10^{-7}$.
Table 12: 1%FPR TPR↓ of watermark stolen on datasets Dolly CW and MMW BookReports.
Method Dolly CW MMW BookReports
KGW 0.45 0.59
KGW (Diff. key) 0.10 0.16
Unigram 0.22 0.04
Unigram (Diff. key) 0.11 0.15
Ours (CD score) 0.13 0.16
Ours (CI score) 0.55 0.56
Ours (CI score, FT w/ $l_{B}$ 5k steps) 0.19 0.12
We conduct a spoof attack (Jovanović et al., 2024) with 2,000 queries per method to evaluate resistance against watermark stolen (lower TPR is better) in Table 12. Our method is more vulnerable with CI scoring (0.55 and 0.56 for the two datasets), but FT w/ reduces the TPR to 0.19 and 0.12, which is comparable to using different keys in KGW and Unigram.
Table 13: Overall performance of our finetune model with logtis bias.
Method 1%FPR TPR↑ PPL↓ BLEU↑ pass@1↑
CL SS CP PA Qlt MT CG
Ours 0.99 0.97 0.98 0.75 7.73 31.06 34.00
Ours (FT) 1.00 0.98 0.99 0.53 7.28 30.94 32.00
Meanwhile, as shown in Table 13, our FT model shows performance comparable to the original checkpoint. Notably, the PPL improves from 7.73 to 7.28 but the TPR drops from 0.75 to 0.53 in the PA case.
B.4 $\gamma$ of Red-Green Split
Table 14: Statistics of $\gamma$ on different corpus.
Hugging Face Dataset Mean ± Std
xsum 0.38 ± 0.11
StackOverflow-QA-C-Language-40k 0.43 ± 0.11
ML-ArXiv-Papers 0.40 ± 0.13
mimic-cxr-dataset 0.40 ± 0.12
finance-alpaca 0.41 ± 0.11
Different from KGW, the ratio of green tokens in the red-green partition $\gamma$ is not deterministic, but train by the watermark encoder, which could be vary with different context. We show the statistics of $\gamma$ in Table 14, and find that our method achieves an average $\gamma$ of 0.4 across datasets with diverse topics.
B.5 Token Distribution Bias
Figure 11: Token distribution and KD divergence (lower the better) between non-watermark/watermark sentences of our method in different settings as well as the SOTA competitors.
We further examine the effect of watermark strength $\delta$ on token distribution, as depicted in Fig. 11. With the default setting of $\delta=1.25$, our model introduces less token bias compared to Unigram and SIR, while achieving superior robustness, as shown in Table 6. However, although our method outperforms KGW in robustness, it introduces a greater token bias. We argue that this highlights an inherent trade-off between robustness and token bias. Reducing $\delta$ to 0.75 reduces the KL divergence by 50% compared to the default setting, but comes at the expense of reduced robustness.
Appendix C Training and Evaluation Details
C.1 Model Training Details
@@
Learning rate 1e-4
Batch size 8
Training step 35k
@@
Training Strategy. To ensure stability and promote convergence in our end-to-end model, we adopt two training strategies: the Multiple-Gradient Descent Algorithm (Huo et al., 2024; Désidéri, 2012) and Curriculum Learning (Bengio et al., 2009).
  1. 1.
    Multiple-Gradient Descent Algorithm. The detection loss $\mathcal{L}_{\text{dec}}$ and the semantic loss $\mathcal{L}_{\text{sem}}$ are inherently conflicting: reducing one often increases the other (Huo et al., 2024). We apply MGDA to resolve this multi‐objective optimization, which is proven to converge to a Pareto stationary solution under mild conditions (Désidéri, 2012).
  2. 2.
    Curriculum Learning. Training is structured into three progressive stages to move from easy to hard:
    • Stage 1 ($<10k$ steps): optimize only the detection objective.
    • Stage 2 ($10k–20k$ steps): optimize both detection and semantic objectives jointly.
    • Stage 3 ($>20k$ steps): activate the noise layer for online paraphrasing together with the two objectives.
    This curriculum schedule enhances training stability and accelerates convergence.
C.2 Training Stability
Figure 12: Training history of our model with three different seeds (63, 64, and 65).
To assess training stability, we train our end-to-end model from scratch using three different seeds (5k steps each) and show the loss history and performance metrics in Fig. 12. We find that our models with different seeds converge in terms of detection loss and accuracy as the training process continues.
C.3 Editing Configuration
@@
Paraphrasing Paraphrasing the text while maintaining the original meaning with the deep paraphrase model Dipper(Krishna et al., 2023). Lex diversity: 60
@@
C.4 Examples for online text modification
@@
Copy-paste Please randomly copy the below given text and paste to the provided human-written text to form the new text: Climate change is one of the most pressing environmental issues of our time, posing significant threats to biodiversity worldwide. Global warming is a critical issue that affects various aspects of our planet. Climate change is one of the most pressing environmental issues of our time, posing significant threats to biodiversity worldwide. It leads to melting ice caps, rising sea levels, and more frequent natural disasters. These changes have profound impacts on both human societies and wildlife. …
@@
Appendix D Comparison with Post-Generation Watermarking Methods
Appendix E Watermark Efficiency
LLM Setting Generation Detection
Time↓ (s) Memory↓ (GB) Time↓ (s) Memory↓ (GB)
OPT-1.3B w/o watermark 2.557 5.900 0.003 0.008
OPT-1.3B w/ watermark 2.769 5.900 0.003 0.008
Llama2-7B w/o watermark 5.204 15.793 0.005 0.008
Llama2-7B w/ watermark 8.362 15.793 0.005 0.008
In Table 18, we assess the computational time overhead and GPU memory usage of our method. For OPT-1.3B, the lightweight encoder design results in only an 8.3% increase in generation time for watermarked text, with no change in maximum GPU memory usage since the encoder is invoked after each token generation. For Llama2-7B, our method increases the generation time by 60.6%, mainly due to the embedding transformation from Llama2-7B to OPT-1.3B, as the tokenizer cannot be accelerated by the GPU and is called at each step. The time overhead for watermark embedding can be mitigated through parallel tokenization, reducing the time complexity by up to $1/k$. In terms of watermark detection, our decoder operates efficiently, requiring only negligible time and memory consumption.
Table 19: Detection time overhead comparison on a single A6000.
Method Detection Time↓ (s)
KGW 0.3
EXP-Edit 80
Unbiased 3.4
Ours 0.005
Moreover, our watermark detection is extremely efficient. As shown in Table 19, with Llama2-7B using a single NVIDIA A6000 GPU, our method requires only 0.005 seconds per watermarked sample (200 tokens), compared to 0.3 seconds for KGW, 3.4 seconds for Unbiased, and 80 seconds for EXP-Edit. Overall, our method is 16,000 times faster than EXP-Edit and 680 times faster than Unbiased, making it highly suitable for scalable watermarking systems.